OpenStudy (anonymous):

@johnweldon1993 this one confuses me A motorboat travels 198 mi in 3 hours going upstream and 416 mi in 4 hours going downstream. What is the rate of the boat in still water and what is the rate of the current?

OpenStudy (anonymous):

if we assume the speed of boat be x and that of the current be y mi/h respectively moving downstream(that is with the current ) relative speed is (x+y) hence 4(x+y)=416 ....(x+y)=104 moving upstream(that is against the current ) relative speed is (x-y) and 3(x-y)=198 ....(x-y)=66 solving for x and y we get the answers

OpenStudy (johnweldon1993):

Thank you @matricked , my wifi really does not like OS....You're in good hands @Mpinkham

OpenStudy (anonymous):

adding the equations we get 2x=170 hence x=85 mi/h putting the value of x in any eq we have y=19

OpenStudy (anonymous):

welcome dear.... even most of the times i also have hardware probs

OpenStudy (anonymous):

so then its 85 and 19??

OpenStudy (johnweldon1993):

That would be correct Check them :D 3(85 - 19) = 198 4(85 + 19) = 416 True?

OpenStudy (anonymous):

for boat its 85 mi/h and for current its 19mi/h

OpenStudy (anonymous):

it is lol